# LARES: Latent Reasoning for Sequential Recommendation

### v1

@@ 2.3.2. Step-Level Alignment
Given a historical interaction sequence $S_{u}$ and a sampled iteration number $k$, LARES generates a sequence of latent thought representations $\bm{T}_{u}=[\bm{t}_{1},\dots,\bm{t}_{k}]$. Ideally, these latent thoughts are progressively refined to converge toward the true user preference distribution. While the model is expected to produce increasingly accurate latent representations as reasoning advances, intermediate states may occasionally diverge from the desired trajectory, leading to suboptimal or counterproductive reasoning steps. To address this issue, we introduce step-level alignment to enforce coherence between intermediate states and the final output. Specifically, we uniformly sample an intermediate step $b$ from $\{1,\dots,k-1\}$ and optimize the alignment between $\bm{o}_{b}$ and $\bm{o}_{k}$ using an InfoNCE loss:
@@
where $\operatorname{Unif}$ denotes the uniform distribution, $s(\cdot,\cdot)$ is a similarity metric (e.g., cosine or dot product), $\mathcal{B}_{i},i\in\{b,k\}$ contains the $i$th step latent reasoning representations of all instances in the same batch, and $\tau$ is a temperature coefficient controlling the distribution sharpness.
@@ 2.4. Reinforcement Post-training for Thinking Exploration
During the self-supervised pre-training stage, the model acquires latent reasoning patterns for sequential recommendation tasks. However, this stage suffers from limited thinking exploration due to the absence of supervisory signals that guide the model in distinguishing between ‘good’ and ‘not good’ reasoning steps. Consequently, the model’s exploratory potential remains underutilized. To mitigate this limitation, we introduce a reinforcement learning-based post-training approach that enhances the model’s reasoning capabilities through learning from experiences of high-quality reasoning trajectories. The subsequent sections introduce the reinforcement learning algorithm, reward design, and data selection startegy.
@@ 2.4.1. Reinforcement Learning Algorithm
To balance performance and computational cost, we employ the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024) during reinforcement post-training. For each input (i.e., user interaction sequences $S_{u}$ in our case), GRPO samples a group of $G$ rollouts $\{y_{1},\dots,y_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ where $G$ denotes the group size. The current policy $\pi_{\theta}$ (i.e., the base model) is then updated by maximizing a surrogate reward, regularized by the KL divergence from a reference policy $\pi_{\theta_{\text{ref}}}$ (i.e., the initial pre-trained model). The formulation of GRPO in our scenario is partially different from the original setting. Specifically, the output $y_{i}$ differs from conventional LLM outputs (i.e., response tokens). Instead, it represents recommended items conditioned on latent thoughts at different steps, denoted as $t^{i}=\{t_{j}^{i}\}_{j=1}^{k}$. To align with sequential recommendation objectives, we assume the target item (rather than the highest-probability candidate) is recommended at each step, thereby explicitly maximizing the target item likelihood. The rollout probability is accordingly defined as the joint probabilities of recommending target items for all latent thinking states: $\pi(y_{i}|S_{u})=\prod_{j=1}^{k}\pi(v_{n}|S_{u},t^{i}_{j})$ where $v_{n}$ denotes the target item. During training, we perform one-step updates on rollouts, maintaining $\pi_{\theta}=\pi_{\theta_{\text{old}}}$. Overall, the GRPO objective is formally expressed as:
(14) $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{S_{u}\sim D,\{t^{i}\}_{i=1}^{G}\sim\pi_{\theta}}\left[\frac{1}{G}\sum_{i=1}^{G}\left(A_{i}-\beta\ \mathbb{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{{\text{ref}}}\right)\right)\right],$
(15) $\mathbb{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{{\text{ref}}}\right)=\prod_{j=1}^{k}\frac{\pi_{{\text{ref}}}(v_{n}|S_{u},t^{i}_{j})}{\pi_{\theta}(v_{n}|S_{u},t^{i}_{j})}-\sum_{j=1}^{k}\log\frac{\pi_{{\text{ref}}}(v_{n}|S_{u},t^{i}_{j})}{\pi_{\theta}(v_{n}|S_{u},t^{i}_{j})}-1.$
@@ 2.4.2. Reward Design
To maintain strict alignment with recommendation objectives, we directly employ standard recommendation metrics (namely NDCG@K and Recall@K) as reward signals. The reward of each rollout is determined by the recommendation results of the last reasoning step and shared by all previous latent reasoning steps. For $G$ rollout trajectories $\{t^{1},\dots,t^{G}\}$, we first obtain the last latent thought states corresponding to the last item as the final user representations $\{\bm{p}_{1},\dots,\bm{p}_{G}\}$ and obtain the recommendation probability distribution for each rollout by computing the similarity between $\bm{p}_{i}$ and item embedding table $\bm{E}$, denoted as $P_{i}=\operatorname{softmax}(\bm{p}_{i}\bm{E}^{T})\in\mathbb{R}^{N},i\in\{1,\dots,G\}$. Then we can obtain the item ranking list based on the probability distribution for every rollout, denoted as $\operatorname{IL}_{i}=\operatorname{argsort}P_{i}$. The reward function is formally written as:
(16) $r_{i}=m(v_{n},\operatorname{IL}_{i}),\quad m\in\{\text{NDCG}@K,\text{Recall}@K\}.$
@@ 2.4.3. Data Selection
To improve the training efficiency of Reinforcement Post-Training (RPT), we propose a data selection strategy that filters out hard training samples to reduce the size of training datasets. Specifically, we exclude instances where the pre-trained model fails to rank the target item within the top-100 candidates across three independent inference trials. This selection criterion is empirically justified, as we observe that such challenging instances contribute minimal reward signals during training and consequently have negligible impact on final model performance.
@@ 2.5. Time Complexity
As described in Section 2.2, the LARES framework comprises two key components: a pre-block and a core-block. To illustrate the computational complexity, we adopt the transformer architecture as a representative example due to LARES’s compatibility with diverse model designs. The primary computational overhead in each transformer layer stems from the Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN) modules, with complexities of $\mathcal{O}(N^{2}d+Nd^{2})$ and $\mathcal{O}(Nd^{2})$ respectively, where $N$ represents the sequence length and $d$ denotes the hidden dimension. Let $L_{1}$ and $L_{2}$ denote the number of transformer layers in the pre-block and core-block, respectively. Consequently, their computational complexities can be expressed as $\mathcal{O}(L_{1}(N^{2}d+Nd^{2}))$ and $\mathcal{O}(L_{2}(N^{2}d+Nd^{2}))$. For a reasoning process involving $K$ iterative steps, the overall time complexity of LARES scales as $\mathcal{O}((L_{1}+L_{2}K)(N^{2}d+Nd^{2}))$.
@@ 3.1.2. Baseline Models
To conduct a comprehensive evaluation of LARES’s performance, we compare it with multiple sequential recommendation (SR) baselines, including both reasoning-based and non-reasoning approaches:
(1) Non-Reasoning SR methods:
  • GRU4Rec (Hidasi et al., 2016a) employs GRUs to capture user behavior patterns.
  • SASRec (Kang and McAuley, 2018) is a transformer-based model utilizing unidirectional multi-head self-attention to encode user interaction sequences.
  • BERT4Rec (Sun et al., 2019) is a bidirectional self-attentive model that employs masked prediction for sequence modeling.
  • FMLP-Rec (Zhou et al., 2022) replaces traditional self-attention with filter-enhanced MLPs to improve behavior modeling.
  • BSARec (Shin et al., 2024) leverages Fourier transforms to capture both high- and low-frequency information in user behavior sequences.
  • CL4SRec (Xie et al., 2022) first introduces contrastive learning for sequential recommendation through three augmentation strategies: item masking, reordering, and cropping.
  • DuoRec (Qiu et al., 2022) combines unsupervised model-level dropout augmentation with hard positive sample selection.
(2) Reasoning-based SR methods:
  • ERL (Tang et al., 2025) enhances sequential recommendations by aggregating multi-step implicit reasoning states.
  • PRL (Tang et al., 2025) improves reasoning capabilities for sequential recommendation through contrastive learning with noise-disturbed positive views and temperature annealing.
  • PRL++ extends PRL by incorporating DuoRec’s sampling strategy.
@@ 3.1.4. Implementation
We implement LARES and all baseline models using PyTorch. To ensure a fair comparison, the batch size, embedding size and hidden size are set to 1024, 64 and 256, respectively. All models are optimized using the AdamW optimizer with a learning rate of 0.001. For self-attentive models, the number of attention heads is fixed at 2. For LARES, the number of layers for both the pre-block and core-block is selected from $\{1,2\}$, while the mean reasoning step $\bar{k}$ is chosen from $\{3,4,6\}$. The hyperparameters for latent reasoning state initialization $\sigma_{1}$ and $\omega$ are set to 1 and 3, respectively. In SPT, we use a learning rate of 0.001 and a dropout rate of 0.5. The coefficients for trajectory-level alignment $\alpha$ and step-level alignment $\gamma$ are tuned within $\{0.1,0.2,0.3\}$ and $\{0.1,0.3,0.5,0.7\}$, respectively. In RPT, the learning rate is selected from $\{0.0005,0.0003,0.0001\}$, and $\beta$ is tuned from $\{0.5,1.0\}$. The rollout number $G$ is fixed at 4. The reward function is selected from $\{\operatorname{Recall@5},\operatorname{Recall@10}\}$. The non-reasoning SR baselines are implemented using RecBole (Zhao et al., 2021, 2022), an open-source recommendation library, while reasoning-based SR models are reproduced from their official source codes. To mitigate overfitting, we employ early stopping, terminating training if NDCG@10 on the validation set shows no improvement for 10 consecutive epochs.
@@ 3.2. Overall Performance
  • Our proposed framework LARES outperforms all baselines, including both non-reasoning and reasoning models, by a large margin across all evaluation metrics on four datasets. This consistent superiority underscores its advanced reasoning capabilities for sequential recommendation tasks. LARES enables multi-step latent reasoning through core-block iteration at a arbitrary depth, allowing flexible computation scaling by increasing the computational density of parameters. To fully exploit its latent reasoning potential, we design two training stages, self-supervised pre-training to instill latent reasoning patterns tailored for sequential recommendation tasks and reinforcement post-training to further stimulate its reasoning capabilities by encouraging thinking exploration. These results also validate the effectiveness of recurrent-depth latent reasoning for sequential recommendation tasks.
@@ 3.3. Ablation Studies
Figure 2. Performance comparison of LARES across different backbone architectures on Instrument and Scientific. ‘Base’ indicates the original backbone model. ‘LR+SPT’ refers to the backbone enhanced with our proposed latent reasoning module and pre-training. ‘LARES’ denotes the full implementation incorporating all proposed components.
@@ 3.4.1. Compatibility with Different Backbones
To evaluate the compatibility of LARES with different architectures, we conduct experiments using three advanced backbones in addition to SASRec (namely FMLP-Rec, BSARec, and TedRec (Xu et al., 2024)). The first two are conventional ID-based sequential recommendation models, whereas TedRec is a text-enhanced model that incorporates both ID embeddings and text embeddings derived from item metadata using a pretrained language model (i.e., Sentence-T5 in our case). Figure 2 presents the performance comparison of three model variants: (1) the backbone (Base), (2) the backbone enhanced with latent reasoning through pre-training (LR+SPT), and (3) the complete framework (LARES). The experimental results demonstrate that LARES consistently improves all four backbones, achieving performance gains of 15.3% and 21.4% in NDCG@10 for FMLP-Rec on the Instrument and Scientific datasets, respectively. For BSARec, LARES yields over 10% improvement across all metrics on both datasets. Notably, even the state-of-the-art TedRec model benefits from our framework, showing average improvements of 10.9% and 7.9% on the respective datasets. These results substantiate the great compatibility of our latent reasoning paradigm. Furthermore, we observe a consistent performance ranking across all metrics and datasets: Base ¡ LR+SPT ¡ LARES. This hierarchy validates the effectiveness of our proposed two-stage training strategy.
@@ 3.4.2. Influence of Reasoning Steps on Recommendation Performance
To analyze the influence of reasoning step, we vary the number of reasoning steps $\bar{k}$ in $\{1,2,3,4,6,8\}$ and evaluate the model on the Instrument and Scientific datasets. The results in Figure 3 reveal a consistent trend: performance initially improves with more reasoning steps but declines after reaching an optimal point, which is in line with the findings in ReaRec. This pattern suggests that moderate reasoning step enhances model performance by strengthening representation power through additional computation. However, excessively large $\bar{k}$ leads to performance degradation, likely because simple user interaction sequences do not require intensive reasoning-a phenomenon analogous to “overthinking” in NLP . The optimal reasoning steps are 3 for Instrument and 2 for Scientific, highlighting the importance of selecting an appropriate reasoning depth for optimal performance.

### v2

@@ 2.3.2. Step-Level Alignment
Given a historical interaction sequence $S_{u}$ and a sampled iteration number $k$, LARES generates a sequence of latent thought representations $\bm{T}_{u}=[\bm{t}_{1},\dots,\bm{t}_{k}]$. Ideally, these latent thoughts are progressively refined to converge toward the true user preference distribution. While the model is expected to produce increasingly accurate latent representations as reasoning advances, intermediate states may occasionally diverge from the desired trajectory, leading to suboptimal or counterproductive reasoning steps. To address this issue, we introduce step-level alignment to enforce coherence between intermediate states and the final output. Specifically, we uniformly sample an intermediate step $b$ from $\{1,\dots,k-1\}$ and optimize the alignment between $\bm{t}_{b}$ and $\bm{t}_{k}$ using an InfoNCE loss:
@@
where $\operatorname{Unif}$ denotes the uniform distribution and $\mathcal{B}_{i},i\in\{b,k\}$ contains the $i$-th step latent reasoning representations of all instances in the same batch.
@@ 2.4. Reinforcement Post-Training for Thinking Exploration
After the self-supervised pre-training stage, the model acquires latent reasoning patterns for sequential recommendation tasks. However, this stage suffers from limited thinking exploration due to the absence of supervisory signals that guide the model in distinguishing between “good” and “not good” reasoning steps. Consequently, the model’s exploratory potential remains underutilized. To mitigate this limitation, we introduce a reinforcement learning-based post-training approach that enhances the model’s reasoning capabilities through learning from experiences of high-quality reasoning trajectories. The subsequent sections introduce the reinforcement learning algorithm, reward design, and data selection strategy.
@@ 2.4.1. Reinforcement Learning Algorithm
To balance performance and computational cost, we employ the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024) during reinforcement post-training. For each input (i.e., user interaction sequences $S_{u}$ in our case), GRPO samples a group of rollouts from the old policy $\pi_{\theta_{\text{old}}}$. The current policy $\pi_{\theta}$ (i.e., the base model) is then updated by maximizing a reward function and regularized by the KL divergence from a reference policy $\pi_{\text{ref}}$ (i.e., the initial pre-trained model). The objective of GRPO is formulated as:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{x\sim D,\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta}}$
(16) $\left[\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(P\cdot A_{i},C\cdot A_{i}\right)-\beta\ \mathbb{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{{\text{ref}}}\right)\right)\right],$
(17) $P=\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},C=\operatorname{clip}\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\epsilon,1+\epsilon\right),$
where $A$ is the advantage value, $x$ is the input, $y_{i}$ is the response generated by LLMs, and $\pi_{\theta}(y|x)=\prod_{j=1}^{|y_{i}|}\pi_{\theta}(y_{i,j}|x,y_{i,<j})$ is the generation probability of the response $y_{i}$ under policy $\pi$. However, directly applying Eq. (16) to our task is infeasible because it requires computing the joint probability $\pi(y_{i}|x)$ with discrete tokens, which are absent in our setting. To resolve this issue, we reformulate it as the joint probability of recommending the target item at each reasoning step, denoted as $\pi(y_{i}|x)=\pi(v_{n+1}|S_{u})=\prod_{j=1}^{k}\pi(v_{n+1}|S_{u},t_{i,j})$ where $v_{n+1}$ is the target item, $S_{u}$ denotes the input user sequence, $t_{i,j}$ represents the latent thought representation at $j$-th step of the $i$-th rollout and $k$ is the total number of reasoning steps.
@@ 2.4.2. Reward Design
To maintain strict alignment with recommendation objectives, we directly employ standard recommendation metrics (namely NDCG@k and Recall@k) as reward signals. We take the recommendation results of the last reasoning step and the target label to calculate these metrics. Specifically, for $G$ rollout trajectories, we first take the last latent thought states corresponding to the last item in the input sequence as the final user representations $\{\bm{p}_{1},\dots,\bm{p}_{G}\}$ and then obtain the recommendation probability distribution for each rollout by computing the similarity between $\bm{p}_{i}$ and the item embedding table $\bm{E}$, denoted as $P_{i}=\operatorname{softmax}(\bm{p}_{i}\cdot\bm{E}^{T})\in\mathbb{R}^{N},i\in\{1,\dots,G\}$. Then we can obtain the item ranking list based on the probability distribution for every rollout, denoted as $\operatorname{L}_{i}=\operatorname{argsort}P_{i}$. The reward function is formally written as:
(18) $r_{i}=m(v_{n+1},\operatorname{L}_{i}),\quad m\in\{\text{NDCG}@k,\text{Recall}@k\}.$
@@ 2.4.3. Data Selection
Recent work demonstrates that the difficulty of data is important to the effectiveness of RL (Team et al., 2025). Since labels in sequential recommendation are very sparse, it is important that the data have a balanced difficulty for a limited number of rollouts. Considering this, we propose a data selection strategy that filters out hard training samples. Specifically, we exclude instances where the pre-trained model fails to rank the target item within the top 100 positions across three independent inference trials.
@@ 2.5. Time Complexity
As described in Section 2.2, the LARES framework comprises two key components: a pre-block and a core-block. To illustrate the computational complexity, we adopt the transformer architecture as a representative example due to LARES’s compatibility with diverse model designs. The primary computational overhead in each transformer layer stems from the multi-head self-attention, with complexities of $\mathcal{O}(N^{2}d+Nd^{2})$, where $N$ represents the sequence length and $d$ denotes the model dimension. Let $L_{1}$ and $L_{2}$ denote the number of transformer layers in the pre-block and core-block, respectively. Consequently, their computational complexities can be expressed as $\mathcal{O}(L_{1}(N^{2}d+Nd^{2}))$ and $\mathcal{O}(L_{2}(N^{2}d+Nd^{2}))$. For a reasoning process involving $K$ iterative steps, the overall time complexity of LARES scales as $\mathcal{O}((L_{1}+L_{2}K)(N^{2}d+Nd^{2}))$.
@@ 3.1.2. Baseline Models
To conduct a comprehensive evaluation of LARES’s performance, we compare it with multiple sequential recommendation baselines, including both reasoning-based and non-reasoning approaches: (1) Non-Reasoning methods: GRU4Rec (Hidasi et al., 2016a) employs GRUs to capture user behavior patterns. SASRec (Kang and McAuley, 2018) is a transformer-based model utilizing unidirectional multi-head self-attention to encode user interaction sequences. BERT4Rec (Sun et al., 2019) is a bidirectional self-attentive model that employs masked prediction for sequence modeling. FMLP-Rec (Zhou et al., 2022) replaces traditional self-attention with filter-enhanced MLPs to improve behavior modeling. BSARec (Shin et al., 2024) leverages Fourier transforms to capture both high- and low-frequency information in user behavior sequences. CL4SRec (Xie et al., 2022) first introduces contrastive learning for sequential recommendation through three augmentation strategies: item masking, reordering, and cropping. DuoRec (Qiu et al., 2022) combines unsupervised model-level dropout augmentation with hard positive sample selection. (2) Reasoning-based methods: ERL (Tang et al., 2025) enhances sequential recommendations by aggregating multi-step implicit reasoning states. PRL (Tang et al., 2025) improves reasoning capabilities for sequential recommendation through contrastive learning with noise-disturbed positive views and temperature annealing. PRL++ extends PRL by incorporating DuoRec’s sampling strategy.
@@ 3.1.4. Implementation
We implement LARES and all baseline models using PyTorch. To ensure a fair comparison, the batch size, embedding size and hidden size are set to 1024, 64 and 256, respectively. All models are optimized using the AdamW optimizer with a learning rate of 0.001. For self-attentive models, the number of attention heads is fixed at 2. For LARES, the number of layers for both the pre-block and core-block is selected from $\{1,2\}$, while the mean reasoning step $\bar{k}$ is chosen from $\{3,4,6\}$. The hyperparameter for latent reasoning state initialization $\sigma_{1}$ is set to 1. In SPT, we use a learning rate of 0.001 and a dropout rate of 0.5. The coefficients for trajectory-level alignment $\alpha$ and step-level alignment $\gamma$ are tuned within $\{0.1,0.2,0.3\}$ and $\{0.1,0.3,0.5,0.7\}$, respectively. In RPT, the learning rate is selected from $\{0.0005,0.0003,0.0001\}$, and $\beta$ is tuned from $\{0.5,1.0\}$. The rollout number $G$ is fixed at 4. The reward function is selected from $\{\operatorname{Recall@5},\operatorname{Recall@10}\}$. The non-reasoning SR baselines are implemented using RecBole (Zhao et al., 2021, 2022), an open-source recommendation library, while reasoning-based SR models are reproduced from their official source codes. To mitigate overfitting, we employ early stopping, terminating training if NDCG@10 on the validation set shows no improvement for 10 consecutive epochs.
@@ 3.2. Overall Performance
  • Our proposed framework LARES outperforms all baselines, including both non-reasoning and reasoning models, by a large margin across all evaluation metrics on four datasets. This consistent superiority underscores its advanced reasoning capabilities for sequential recommendation tasks. LARES enables multi-step latent reasoning through core-block iteration at an arbitrary depth, allowing flexible computation scaling by increasing the computational density of parameters. To fully exploit its latent reasoning potential, we design two training stages: self-supervised pre-training to instill latent reasoning patterns tailored for sequential recommendation tasks and reinforcement post-training to further stimulate its reasoning capabilities by encouraging thinking exploration. These results also validate the effectiveness of depth-recurrent latent reasoning for sequential recommendation tasks.
@@ 3.3. Ablation Studies
Figure 2. Performance comparison of LARES across different backbone architectures on Instrument and Scientific. ‘Base’ indicates the original backbone model. ‘LR+SPT’ refers to the backbone enhanced with our proposed latent reasoning module and pretraining. ‘LARES’ denotes the full implementation incorporating all proposed components.
@@ 3.4.1. Compatibility with Different Backbones
To evaluate the compatibility of LARES with different architectures, we conduct experiments using three advanced backbones in addition to SASRec (namely FMLP-Rec, BSARec, and TedRec (Xu et al., 2024)). The first two are conventional ID-based sequential recommendation models, whereas TedRec is a text-enhanced model that incorporates both ID embeddings and text embeddings derived from item metadata using a pretrained language model (i.e., Sentence-T5 in our case). Figure 2 presents the performance comparison of three model variants: (1) the backbone (Base), (2) the backbone enhanced with latent reasoning through pretraining (LR+SPT), and (3) the complete framework (LARES). The experimental results demonstrate that LARES consistently improves all four backbones, achieving performance gains of 15.3% and 21.4% in NDCG@10 for FMLP-Rec on the Instrument and Scientific datasets, respectively. For BSARec, LARES yields over 10% improvement across all metrics on both datasets. Notably, even the superior TedRec model benefits from our framework, showing average improvements of 10.9% and 7.9% on the respective datasets. These results substantiate the great compatibility of our latent reasoning paradigm. Furthermore, we observe a consistent performance ranking across all metrics and datasets: Base ¡ LR+SPT ¡ LARES. This hierarchy validates the effectiveness of our proposed two-stage training strategy.
@@ 3.4.2. Influence of Reasoning Steps on Recommendation Performance
To analyze the influence of reasoning step, we vary the number of reasoning steps $\bar{k}$ in $\{1,2,3,4,6,8\}$ and evaluate the model on the Instrument and Scientific datasets. The results in Figure 3 reveal a consistent trend: performance initially improves with more reasoning steps but declines after reaching an optimal point, which is in line with the findings in ReaRec. This pattern suggests that moderate reasoning step enhances model performance by strengthening representation power through additional computation. However, excessively large $\bar{k}$ leads to performance degradation, likely because simple user interaction sequences do not require intensive reasoning-a phenomenon analogous to “overthinking” in NLP (Su et al., 2025). The optimal reasoning steps are 4 for Instrument and 3 for Scientific, highlighting the importance of selecting an appropriate reasoning depth for optimal performance.